# Linear combinations of latents in generative models: subspaces and beyond

### v7

@@ Appendix G Additional qualitative results
Figure 8: Low p-values are inherited by interpolants The x-axis show the lower Kolmogorov-Smirnov p-value of the two seed latents for each corresponding spherical interpolation in Figure 5, and the y-axis show the p-value of the resulting interpolant. Due to the vast dynamic range of the p-values, the scatter plot is shown in full on the left, and zoomed in on the upper-right quadrant on the right. We note that when any of the two seed latents have a small p-value this largely tend to be inherited by the interpolant. The Pearson correlation coefficient is $0.79$ for the indicator of acceptance (defined as $1$ if the p-value is greater than $1e^{-3}$ and 00 otherwise) and $0.66$ for the $\log_{10}$ p-values.
See Figure 9 for a demonstration of subspaces of latents without the LOL transformation introduced in Equation 3, using a diffusion model (Rombach et al., 2022) and a flow matching model (Esser et al., 2024), respectively. The setup is identical (with the same original latents) as in Figure 12 and Figure 1, respectively, except without applying the proposed (LOL) transformation.
@@
Figure 10: Additional slices of a sports car subspace. The latents $\bm{x}_{1},\dots,\bm{x}_{5}$ (corresponding to images) are converted into basis vectors and used to define a 5-dimensional subspace. The grids show generations from uniform grid points in the subspace coordinate system, where the left and right grids are for the dimensions $\{1,2\}$ and $\{3,4\}$, respectively, centered around the coordinate for $\bm{x}_{1}$. Each coordinate in the subspace correspond to a linear combination of the basis vectors. The flow matching model Stable Diffusion 3 (Esser et al., 2024) is used in this example. See Figure 1 for another slice of the latent subspace.
@@
Figure 11: Interpolation. Shown are generations from equidistant interpolations of latents $\bm{x}_{1}$ and $\bm{x}_{2}$ (in $v\in[0,1]$) using the respective method. The diffusion model Stable Diffusion 2.1 (Rombach et al., 2022) is used in this example.
@@
Figure 12: Low-dimensional subspaces. The latents $\bm{x}_{1},\dots,\bm{x}_{5}$ (corresponding to images) are converted into basis vectors and used to define a 5-dimensional subspace. The grids show generations from uniform grid points in the subspace coordinate system, where the left and right grids are for the dimensions $\{1,3\}$ and $\{2,4\}$, respectively, centered around the coordinate for $\bm{x}_{1}$. Each coordinate in the subspace correspond to a linear combination of the basis vectors. The diffusion model Stable Diffusion 2.1 (Rombach et al., 2022) is used in this example.
@@
Figure 13: Low-dimensional subspaces. The latents $\bm{x}_{1},\dots,\bm{x}_{5}$ (corresponding to images) are converted into basis vectors and used to define a 5-dimensional subspace. The grids show generations from uniform grid points in the subspace coordinate system, where the left and right grids are for the dimensions $\{1,3\}$ and $\{2,4\}$, respectively, centered around the coordinate for $\bm{x}_{1}$. Each coordinate in the subspace correspond to a linear combination of the basis vectors. The diffusion model Stable Diffusion 2.1 (Rombach et al., 2022) is used in this example.
@@
Figure 14: Interpolation during unconditional generation. Shown are generations from equidistant interpolations of latents $\bm{x}_{1}$ and $\bm{x}_{2}$ using the respective method. The latents $\bm{x}_{1}$ and $\bm{x}_{2}$ were obtained from DDIM inversion (Song et al., 2020a) with an empty prompt, and all generations in this example are then carried out with an empty prompt. The larger images show interpolation index 8 for SLERP, NAO and LOL, respectively. The diffusion model Stable Diffusion 2.1 (Rombach et al., 2022) is used in this example.
@@ Appendix H Additional quantitative results and analysis
In Table 1 we show that our method outperforms or maintains the performance of the baselines in terms of FID distances and accuracy, as calculated using a pre-trained classifier following the evaluation methodology of Samuel et al. (2023). For an illustration of centroids and interpolations, see Figure 2 and Figure 16, respectively. The evaluation time of an interpolation path and centroid, shown with one digit of precision, illustrate that the closed-form expressions are significantly faster than NAO. Surprisingly, NAO did not perform as well as spherical interpolation and several other baselines, despite using their implementation, which was outperforming these methods in Samuel et al. (2023). We note one discrepancy is that we report FID distances using (2048) high-level features, while in their work they are using (64) low-level features, which in Seitzer (2020) is recommended against as it does not necessarily correlate with visual quality. In Table 3 we report results with all feature settings. We note that, in our setup, the baselines perform substantially better than reported in Samuel et al. (2023) — including NAO in terms of FID distances using the 64 low-level feature setting (1.30 in our setup vs 6.78 in their setup), and class accuracy during interpolation (62% vs 52%), which we study below.
The number of DDIM inversion steps used for the ImageNet images is not reported in the NAO setup (Samuel et al., 2023), but if we use 1 step instead of 400 (400 is what we use in all our experiments) the class preserving accuracy of NAO for centroid determination, shown in Table 4, resemble theirs more, yielding 58%. As we illustrate in Figure 5 and Figure 15, a sufficient number of inversion steps is critical to acquire latents which not only reconstructs the original image but which can be successfully interpolated. We hypothesise that a much lower setting of DDIM inversions steps were used in their setup than in ours, which would be consistent with their norm-correcting optimization method having a large effect, making otherwise blurry/superimposed images intelligible for the classification model.
@@
LERP $3.92$% $63.4422$ $278.0981$ $2.1761$ $198.8888$ $6e^{-3}$s
SLERP $64.56$% $2.2798$ $4.9546$ $0.1727$ $42.5827$ $9e^{-3}$s
NAO 62.133 333 333 333 3362.1333333333333362.133\,333\,333\,333\,3362.133 333 333 333 33% $1.3001$ $4.1055$ $0.1948$ $46.0230$ $30$s
LOL (ours) 67.386 666 666 666 6667.3866666666666667.386\,666\,666\,666\,6667.386 666 666 666 66% $1.7248$ $3.6204$ $0.1559$ $38.8714$ $6e^{-3}$s
@@
Euclidean 0.285 714 285 714 285 70.28571428571428570.285\,714\,285\,714\,285\,70.285 714 285 714 285 7% 67.718 855 147 659 1667.7188551476591667.718\,855\,147\,659\,1667.718 855 147 659 16 317.418 673 402 618 14317.41867340261814317.418\,673\,402\,618\,14317.418 673 402 618 14 3.682 238 270 823 695 43.68223827082369543.682\,238\,270\,823\,695\,43.682 238 270 823 695 4 310.415 342 373 754 75310.41534237375475310.415\,342\,373\,754\,75310.415 342 373 754 75 $4e^{-4}$s
Standardized Euclidean 44.571 428 571 428 57344.57142857142857344.571\,428\,571\,428\,57344.571 428 571 428 573% 5.920 945 934 429 665.920945934429665.920\,945\,934\,429\,665.920 945 934 429 66 21.026 935 293 432 521.026935293432521.026\,935\,293\,432\,521.026 935 293 432 5 0.422 518 772 177 195 250.422518772177195250.422\,518\,772\,177\,195\,250.422 518 772 177 195 25 88.839 964 830 394 788.839964830394788.839\,964\,830\,394\,788.839 964 830 394 7 $1e^{-3}$s
Mode norm Euclidean 44.571 428 571 428 57344.57142857142857344.571\,428\,571\,428\,57344.571 428 571 428 573% 6.910 416 861 443 1076.9104168614431076.910\,416\,861\,443\,1076.910 416 861 443 107 21.564 254 709 507 08221.56425470950708221.564\,254\,709\,507\,08221.564 254 709 507 082 0.455 027 076 332 581 440.455027076332581440.455\,027\,076\,332\,581\,440.455 027 076 332 581 44 88.377 760 874 495 3888.3777608744953888.377\,760\,874\,495\,3888.377 760 874 495 38 $1e^{-3}$s
NAO $44.000$% 4.160 326 379 731 8564.1603263797318564.160\,326\,379\,731\,8564.160 326 379 731 856 15.551 897 857 700 89415.55189785770089415.551\,897\,857\,700\,89415.551 897 857 700 894 0.465 751 253 666 584 650.465751253666584650.465\,751\,253\,666\,584\,650.465 751 253 666 584 65 92.967 278 310 416 1392.9672783104161392.967\,278\,310\,416\,1392.967 278 310 416 13 $90$s
LOL (ours) 46.285 714 285 714 28646.28571428571428646.285\,714\,285\,714\,28646.285 714 285 714 286% 9.377 513 999 920 7089.3775139999207089.377\,513\,999\,920\,7089.377 513 999 920 708 25.545 423 775 866 01325.54542377586601325.545\,423\,775\,866\,01325.545 423 775 866 013 0.454 830 163 321 727 10.45483016332172710.454\,830\,163\,321\,727\,10.454 830 163 321 727 1 87.713 436 033 613 3387.7134360336133387.713\,436\,033\,613\,3387.713 436 033 613 33 $6e^{-4}$s
Table 3: Full table with class accuracy and FID distances using each setting in Seitzer (2020). Note that FID 64, 192, and 768 is not recommended by Seitzer (2020), as it does not necessarily correlate with visual quality. FID 2048 is based on the final features of InceptionV3 (as in Heusel et al. (2017)), while FID 64, 192 and 768 are based on earlier layers; FID 64 being the first layer.
@@
Euclidean 42.285 714 285 714 2942.2857142857142942.285\,714\,285\,714\,2942.285 714 285 714 29% 25.678 102 466 680 16625.67810246668016625.678\,102\,466\,680\,16625.678 102 466 680 166 59.339 892 162 255 3859.3398921622553859.339\,892\,162\,255\,3859.339 892 162 255 38 1.233 642 013 197 828 31.23364201319782831.233\,642\,013\,197\,828\,31.233 642 013 197 828 3 170.438 344 463 461 02170.43834446346102170.438\,344\,463\,461\,02170.438 344 463 461 02 $4e^{-4}$s
Standardized Euclidean 46.571 428 571 428 5746.5714285714285746.571\,428\,571\,428\,5746.571 428 571 428 57% 3.352 239 999 972 9733.3522399999729733.352\,239\,999\,972\,9733.352 239 999 972 973 16.551 997 754 711 73516.55199775471173516.551\,997\,754\,711\,73516.551 997 754 711 735 1.225 691 029 025 259 81.22569102902525981.225\,691\,029\,025\,259\,81.225 691 029 025 259 8 175.212 685 969 288 83175.21268596928883175.212\,685\,969\,288\,83175.212 685 969 288 83 $1e^{-3}$s
Mode norm Euclidean 50.571 428 571 428 5750.5714285714285750.571\,428\,571\,428\,5750.571 428 571 428 57% 2.077 423 435 315 104 42.07742343531510442.077\,423\,435\,315\,104\,42.077 423 435 315 104 4 8.742 233 873 370 688.742233873370688.742\,233\,873\,370\,688.742 233 873 370 68 1.193 557 663 163 881 31.19355766316388131.193\,557\,663\,163\,881\,31.193 557 663 163 881 3 172.829 414 025 984 3172.8294140259843172.829\,414\,025\,984\,3172.829 414 025 984 3 $1e^{-3}$s
NAO 57.714 285 714 285 7257.7142857142857257.714\,285\,714\,285\,7257.714 285 714 285 72% 3.769 826 574 088 071 43.76982657408807143.769\,826\,574\,088\,071\,43.769 826 574 088 071 4 13.348 910 880 007 6813.3489108800076813.348\,910\,880\,007\,6813.348 910 880 007 68 1.053 665 908 903 870 51.05366590890387051.053\,665\,908\,903\,870\,51.053 665 908 903 870 5 150.454 019 109 047 12150.45401910904712150.454\,019\,109\,047\,12150.454 019 109 047 12 $90$s
LOL (ours) 48.571 428 571 428 5748.5714285714285748.571\,428\,571\,428\,5748.571 428 571 428 57% 2.962 489 894 597 137 42.96248989459713742.962\,489\,894\,597\,137\,42.962 489 894 597 137 4 12.460 956 090 514 72812.46095609051472812.460\,956\,090\,514\,72812.460 956 090 514 728 1.219 060 048 266 758 61.21906004826675861.219\,060\,048\,266\,758\,61.219 060 048 266 758 6 171.201 024 385 016 07171.20102438501607171.201\,024\,385\,016\,07171.201 024 385 016 07 $6e^{-4}$s
@@
Figure 15: Accurate DDIM inversions are critical for interpolation. Shown is the interpolation (center) of Image 1 and Image 2 using each respective method, after a varying number of budgets of DDIM inversion steps (Song et al., 2020a). For each budget setting, the inversion was run from the beginning. We note that although a single step of DDIM inversion yields latents which perfectly reconstructs the original image such latents do not lead to realistic images, but merely visual ”superpositions” of the images being interpolated.
@@ Appendix I Additional ImageNet interpolations and centroids
Figure 16: Interpolation. Shown are generations from equidistant interpolations of latents $\bm{x}_{1}$ and $\bm{x}_{2}$ using the respective method. The latents $\bm{x}_{1}$ and $\bm{x}_{2}$ were obtained from DDIM inversion (Song et al., 2020a) of two random image examples from one of the randomly selected ImageNet classes (”Chesapeake Bay retriever”) described in Section 5. The larger images show interpolation index 6 for SLERP, NAO and LOL, respectively. The diffusion model Stable Diffusion 2.1 (Rombach et al., 2022) is used in this example.
@@
Figure 17: ImageNet centroid determination. The centroid of the latents $\bm{x}_{1}$, $\bm{x}_{2}$, $\bm{x}_{3}$ as determined using the different methods, with the result shown in the respective (right-most) plot. The diffusion model Stable Diffusion 2.1 (Rombach et al., 2022) is used in this example.
@@
Figure 18: Centroid determination of many latents. The centroid of the latents as determined using the different methods, with the result shown in the respective (right-most) plot. Ten and 25 latents are used in the first two and second two examples (rows), respectively, and the plots show a subset of these in the left-most plots. We noted during centroid determination of many latents (such as 10 and 25) that the centroids were often unsatisfactory using all methods; blurry, distorted etc. For applications needing to find meaningful centroids of a large number of latents, future work is needed. The diffusion model Stable Diffusion 2.1 (Rombach et al., 2022) is used in this example.

### v8

@@ Appendix G Additional qualitative results
Figure 8: Low p-values are inherited by interpolants The x-axis show the lower Kolmogorov-Smirnov p-value of the two seed latents for each corresponding spherical interpolation in Figure 5, and the y-axis show the p-value of the resulting interpolant. Due to the vast dynamic range of the p-values, the scatter plot is shown in full on the left, and zoomed in on the upper-right quadrant on the right. We note that when any of the two seed latents have a small p-value this largely tend to be inherited by the interpolant. The Pearson correlation coefficient is $0.79$ for the indicator of acceptance (defined as 11 if the p-value is greater than $1e^{-3}$ and 0 otherwise) and $0.66$ for the $\log_{10}$ p-values.
See Figure 9 for a demonstration of subspaces of latents without the LOL transformation introduced in Equation 3, using a diffusion model (rombach2022high) and a flow matching model (esser2024scaling), respectively. The setup is identical (with the same original latents) as in Figure 12 and Figure 1, respectively, except without applying the proposed (LOL) transformation.
@@
Figure 10: Additional slices of a sports car subspace. The latents $\bm{x}_{1},\dots,\bm{x}_{5}$ (corresponding to images) are converted into basis vectors and used to define a 5-dimensional subspace. The grids show generations from uniform grid points in the subspace coordinate system, where the left and right grids are for the dimensions $\{1,2\}$ and $\{3,4\}$, respectively, centered around the coordinate for $\bm{x}_{1}$. Each coordinate in the subspace correspond to a linear combination of the basis vectors. The flow matching model Stable Diffusion 3 (esser2024scaling) is used in this example. See Figure 1 for another slice of the latent subspace.
@@
Figure 11: Interpolation. Shown are generations from equidistant interpolations of latents $\bm{x}_{1}$ and $\bm{x}_{2}$ (in $v\in[0,1]$) using the respective method. The diffusion model Stable Diffusion 2.1 (rombach2022high) is used in this example.
@@
Figure 12: Low-dimensional subspaces. The latents $\bm{x}_{1},\dots,\bm{x}_{5}$ (corresponding to images) are converted into basis vectors and used to define a 5-dimensional subspace. The grids show generations from uniform grid points in the subspace coordinate system, where the left and right grids are for the dimensions $\{1,3\}$ and $\{2,4\}$, respectively, centered around the coordinate for $\bm{x}_{1}$. Each coordinate in the subspace correspond to a linear combination of the basis vectors. The diffusion model Stable Diffusion 2.1 (rombach2022high) is used in this example.
@@
Figure 13: Low-dimensional subspaces. The latents $\bm{x}_{1},\dots,\bm{x}_{5}$ (corresponding to images) are converted into basis vectors and used to define a 5-dimensional subspace. The grids show generations from uniform grid points in the subspace coordinate system, where the left and right grids are for the dimensions $\{1,3\}$ and $\{2,4\}$, respectively, centered around the coordinate for $\bm{x}_{1}$. Each coordinate in the subspace correspond to a linear combination of the basis vectors. The diffusion model Stable Diffusion 2.1 (rombach2022high) is used in this example.
@@
Figure 14: Interpolation during unconditional generation. Shown are generations from equidistant interpolations of latents $\bm{x}_{1}$ and $\bm{x}_{2}$ using the respective method. The latents $\bm{x}_{1}$ and $\bm{x}_{2}$ were obtained from DDIM inversion (song2020denoising) with an empty prompt, and all generations in this example are then carried out with an empty prompt. The larger images show interpolation index 8 for SLERP, NAO and LOL, respectively. The diffusion model Stable Diffusion 2.1 (rombach2022high) is used in this example.
@@ Appendix H Additional quantitative results and analysis
In Table 1 we show that our method outperforms or maintains the performance of the baselines in terms of FID distances and accuracy, as calculated using a pre-trained classifier following the evaluation methodology of samuel2024norm. For an illustration of centroids and interpolations, see Figure 2 and Figure 16, respectively. The evaluation time of an interpolation path and centroid, shown with one digit of precision, illustrate that the closed-form expressions are significantly faster than NAO. Surprisingly, NAO did not perform as well as spherical interpolation and several other baselines, despite using their implementation, which was outperforming these methods in samuel2024norm. We note one discrepancy is that we report FID distances using (2048) high-level features, while in their work they are using (64) low-level features, which in Seitzer2020FID is recommended against as it does not necessarily correlate with visual quality. In Table 3 we report results with all feature settings. We note that, in our setup, the baselines perform substantially better than reported in samuel2024norm — including NAO in terms of FID distances using the 64 low-level feature setting (1.30 in our setup vs 6.78 in their setup), and class accuracy during interpolation (62% vs 52%), which we study below.
The number of DDIM inversion steps used for the ImageNet images is not reported in the NAO setup (samuel2024norm), but if we use 1 step instead of 400 (400 is what we use in all our experiments) the class preserving accuracy of NAO for centroid determination, shown in Table 4, resemble theirs more, yielding 58%. As we illustrate in Figure 5 and Figure 15, a sufficient number of inversion steps is critical to acquire latents which not only reconstructs the original image but which can be successfully interpolated. We hypothesise that a much lower setting of DDIM inversions steps were used in their setup than in ours, which would be consistent with their norm-correcting optimization method having a large effect, making otherwise blurry/superimposed images intelligible for the classification model.
@@
LERP 3.923.92% $63.4422$ $278.0981$ $2.1761$ $198.8888$ $6e^{-3}$s
SLERP 64.5664.56% $2.2798$ $4.9546$ $0.1727$ $42.5827$ $9e^{-3}$s
NAO 62.133 333 333 333 3362.133\,333\,333\,333\,33% $1.3001$ $4.1055$ $0.1948$ $46.0230$ $30$s
LOL (ours) 67.386 666 666 666 6667.386\,666\,666\,666\,66% $1.7248$ $3.6204$ $0.1559$ $38.8714$ $6e^{-3}$s
@@
Euclidean 0.285 714 285 714 285 70.285\,714\,285\,714\,285\,7% 67.718 855 147 659 1667.718\,855\,147\,659\,16 317.418 673 402 618 14317.418\,673\,402\,618\,14 3.682 238 270 823 695 43.682\,238\,270\,823\,695\,4 310.415 342 373 754 75310.415\,342\,373\,754\,75 $4e^{-4}$s
Standardized Euclidean 44.571 428 571 428 57344.571\,428\,571\,428\,573% 5.920 945 934 429 665.920\,945\,934\,429\,66 21.026 935 293 432 521.026\,935\,293\,432\,5 0.422 518 772 177 195 250.422\,518\,772\,177\,195\,25 88.839 964 830 394 788.839\,964\,830\,394\,7 $1e^{-3}$s
Mode norm Euclidean 44.571 428 571 428 57344.571\,428\,571\,428\,573% 6.910 416 861 443 1076.910\,416\,861\,443\,107 21.564 254 709 507 08221.564\,254\,709\,507\,082 0.455 027 076 332 581 440.455\,027\,076\,332\,581\,44 88.377 760 874 495 3888.377\,760\,874\,495\,38 $1e^{-3}$s
NAO 44.00044.000% 4.160 326 379 731 8564.160\,326\,379\,731\,856 15.551 897 857 700 89415.551\,897\,857\,700\,894 0.465 751 253 666 584 650.465\,751\,253\,666\,584\,65 92.967 278 310 416 1392.967\,278\,310\,416\,13 $90$s
LOL (ours) 46.285 714 285 714 28646.285\,714\,285\,714\,286% 9.377 513 999 920 7089.377\,513\,999\,920\,708 25.545 423 775 866 01325.545\,423\,775\,866\,013 0.454 830 163 321 727 10.454\,830\,163\,321\,727\,1 87.713 436 033 613 3387.713\,436\,033\,613\,33 $6e^{-4}$s
Table 3: Full table with class accuracy and FID distances using each setting in Seitzer2020FID. Note that FID 64, 192, and 768 is not recommended by Seitzer2020FID, as it does not necessarily correlate with visual quality. FID 2048 is based on the final features of InceptionV3 (as in heusel2017gans), while FID 64, 192 and 768 are based on earlier layers; FID 64 being the first layer.
@@
Euclidean 42.285 714 285 714 2942.285\,714\,285\,714\,29% 25.678 102 466 680 16625.678\,102\,466\,680\,166 59.339 892 162 255 3859.339\,892\,162\,255\,38 1.233 642 013 197 828 31.233\,642\,013\,197\,828\,3 170.438 344 463 461 02170.438\,344\,463\,461\,02 $4e^{-4}$s
Standardized Euclidean 46.571 428 571 428 5746.571\,428\,571\,428\,57% 3.352 239 999 972 9733.352\,239\,999\,972\,973 16.551 997 754 711 73516.551\,997\,754\,711\,735 1.225 691 029 025 259 81.225\,691\,029\,025\,259\,8 175.212 685 969 288 83175.212\,685\,969\,288\,83 $1e^{-3}$s
Mode norm Euclidean 50.571 428 571 428 5750.571\,428\,571\,428\,57% 2.077 423 435 315 104 42.077\,423\,435\,315\,104\,4 8.742 233 873 370 688.742\,233\,873\,370\,68 1.193 557 663 163 881 31.193\,557\,663\,163\,881\,3 172.829 414 025 984 3172.829\,414\,025\,984\,3 $1e^{-3}$s
NAO 57.714 285 714 285 7257.714\,285\,714\,285\,72% 3.769 826 574 088 071 43.769\,826\,574\,088\,071\,4 13.348 910 880 007 6813.348\,910\,880\,007\,68 1.053 665 908 903 870 51.053\,665\,908\,903\,870\,5 150.454 019 109 047 12150.454\,019\,109\,047\,12 $90$s
LOL (ours) 48.571 428 571 428 5748.571\,428\,571\,428\,57% 2.962 489 894 597 137 42.962\,489\,894\,597\,137\,4 12.460 956 090 514 72812.460\,956\,090\,514\,728 1.219 060 048 266 758 61.219\,060\,048\,266\,758\,6 171.201 024 385 016 07171.201\,024\,385\,016\,07 $6e^{-4}$s
@@
Figure 15: Accurate DDIM inversions are critical for interpolation. Shown is the interpolation (center) of Image 1 and Image 2 using each respective method, after a varying number of budgets of DDIM inversion steps (song2020denoising). For each budget setting, the inversion was run from the beginning. We note that although a single step of DDIM inversion yields latents which perfectly reconstructs the original image such latents do not lead to realistic images, but merely visual ”superpositions” of the images being interpolated.
@@ Appendix I Additional ImageNet interpolations and centroids
Figure 16: Interpolation. Shown are generations from equidistant interpolations of latents $\bm{x}_{1}$ and $\bm{x}_{2}$ using the respective method. The latents $\bm{x}_{1}$ and $\bm{x}_{2}$ were obtained from DDIM inversion (song2020denoising) of two random image examples from one of the randomly selected ImageNet classes (”Chesapeake Bay retriever”) described in Section 5. The larger images show interpolation index 6 for SLERP, NAO and LOL, respectively. The diffusion model Stable Diffusion 2.1 (rombach2022high) is used in this example.
@@
Figure 17: ImageNet centroid determination. The centroid of the latents $\bm{x}_{1}$, $\bm{x}_{2}$, $\bm{x}_{3}$ as determined using the different methods, with the result shown in the respective (right-most) plot. The diffusion model Stable Diffusion 2.1 (rombach2022high) is used in this example.
@@
Figure 18: Centroid determination of many latents. The centroid of the latents as determined using the different methods, with the result shown in the respective (right-most) plot. Ten and 25 latents are used in the first two and second two examples (rows), respectively, and the plots show a subset of these in the left-most plots. We noted during centroid determination of many latents (such as 10 and 25) that the centroids were often unsatisfactory using all methods; blurry, distorted etc. For applications needing to find meaningful centroids of a large number of latents, future work is needed. The diffusion model Stable Diffusion 2.1 (rombach2022high) is used in this example.